OpenStudy (anonymous):

What is the range of the inverse of relation {(1, 7), (2, -4), (5, 6), (2, 8)}? A. {1, 2, 5} B. {-4, 6, 7, 8} C. {1, 5} D. {-4, 7, 8}

OpenStudy (anonymous):

\[\left( 7,1 \right), \left( -4,2 \right)\left( 6,5 \right)\left( 8,2 \right)\]

OpenStudy (anonymous):

this is the inverse ^^^^^^^^^^^^

OpenStudy (anonymous):

\[RANGE=\left\{ 1,2,5 \right\}\]

OpenStudy (anonymous):

C) (1,5) the reason is your relation is pairs of number. consider the first number as x, and the second is y . your original function has x range from: 1,2,2,5 means the range is 1,5. from now on. you can say that the inverse is x is replace by y. so, in inverse function, you have y range from 1,5. and because it is inverse, the range does not depend on x value any more, but on y. so the answer is 1,5. I just want you to understand why the result is that and you can calculate by yourself in the future. Hope this help. @ some...someone. I'm sorry if I bother you. I just want to help the friend a little bit when i can.

OpenStudy (anonymous):

your the best @Hoa

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No, the range of the inverse are the y-values. which is (1,2,5)

OpenStudy (anonymous):

@some...someone. you are right. 100% right. I don't put 2 in the middle, because it is there already when the range go from1 to 5 , it includes 2